# ToDo (#1): #2

### v3

@@ 2.1.3 Electrical Design
Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig. 2.2. The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range. This system interfaces with the EPICS PC control system which is documented in detail in Ref. [weigel_epics_2023].
The filament is heated by an TDK Lambda (8.4 V, 300 A max) power supply. In addition to this power supply which is used to heat the filament, a second power supply (Matsusada REK) is connected to the filament and to the source chamber to maintain a potential difference of 80 - 140 V between them in order to facilitate electron discharge of up to 24 A max (typical during our tests was 5-10 A). The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape. Power is transferred to the platform via two isolation transformers of 2 kVA each. A schematic of the wiring of the ion source is illustrated in Fig. 2.3.
The ion source (platform) is lifted to high voltage using a Matsusada AU 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU 7.5 mA, 20 kV power supplies of either negative or positive polarity.
@@
Figure 2.3: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
@@ 2.1.4 Beam Dynamics
We simulate beam formation and shaping in the extraction system with the IBSimu package [kalvasIBSIMUThreedimensionalSimulation2010a], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process. IBSimu also contains a 1D plasma sheath model which is a good approximation for the plasma-beam interface at the source aperture. Our extraction simulations of the MIST-1 ion source followed by a WARP [vayNovelMethodsParticleInCell2012] simulation of the LEBT yielded excellent agreement with measurements [winklehner_first_2018, winklehner_new_2022]. Example trajectories of the new MIST-1 extraction system and RFQ matching LEBT are shown in Fig. 2.1.
An extraction system simulation of the MIST-2 system is shown in Fig. 2.4. The simulated normalized RMS emittance for the new MIST-2 extraction system using an 8 mm diameter plasma aperture is 0.077 $\pi$-mm-mrad. Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [winklehner_first_2018].
Figure 2.4: IBSimu Simulation of MIST-2 extraction. Left: trajectories, Right: x-x’ phase space. Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
@@ 2.1.5 Performance
2.1.5.1 Goals & Measurement Methods
@@
  • High current: $\approx$10 mA of H+${}_{2}^{+}$
  • High purity: H+${}_{2}^{+}$ fraction $>$ 80%
  • Low emittance: $<$ 0.1 $\pi$-mm-mrad (RMS, normalized)
@@
2.1.5.2 Optimization of Ion Source Parameters
@@
H2 input flow optimization. Three series of optimizations on the input gas flow rate have been performed. The three series differ in how the filament discharge voltage is managed. In the first series, the PID loop which stabilizes discharge was disabled allowing the discharge current to rise to the source steady-state value and then stabilized at that value by reengaging the PID. In the second series, the discharge current was kept fixed at 4 A. In the third series, the discharge current was reduced to keep the total extracted current fixed at 1.25 mA. Under each of these three conditions, the gas input rate was varied from 0.25 - 1.25 SCCM. This study demonstrates a general trend of greater H+${}_{2}^{+}$ production at higher discharge currents and, as theoretically expected, a higher fraction of H+${}_{3}^{+}$ with higher flow rates.
@@
Discharge current optimization. One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A. No strong dependence on current was observed.
2.1.5.3 Latest Optimized Results
To date, only an incomplete optimization of the ion source has been performed. As such, parameters which simultaneously meet all the target goals have not yet been fully identified. The ion source’s peak performance to date varies, therefore, depending on which of the performance metrics is optimized. Optimization has been performed for maximum total current density, maximum H+${}_{2}^{+}$ purity, and a hybrid option optimizing for the maximum H+${}_{2}^{+}$ current density with the constraint that H+${}_{2}^{+}$ be dominant.
Highest total current density: 41 mA/cm2 with 31% H+${}_{2}^{+}$ purity.
Highest H+${}_{2}^{+}$ purity: 76% purity with 11.4 mA/cm2 total current density.
Highest H+${}_{2}^{+}$ current density with H+${}_{2}^{+}$ dominant: 10.4 mA/cm2 with 62% purity [^2] Note that this is the current density of $\mathrm{H}_{2}^{+}$ alone..
With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity. With an 8 mm aperture this corresponds to 4.4 mA. This is a world-leading result and is only about a factor of 2 short of the target.
The main bottleneck for an increased beam current was the extraction system, which was designed for initial testing and commissioning. We have recently replaced it with the version presented in this chapter. Our high-fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current. Measurements are planned for summer 2024.
2.1.5.4 Magnet Study
@@
We aim to maximize the $\mathrm{H}_{2}^{+}$ fraction of the MIST-1 ion source. Within the plasma, atomic processes include ionization, dissociation, and recombination, leading to protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$ in the beam. Magnets with a higher field strength yield higher plasma density and longer confinement of the ions in the plasma, which can alter the rates of these processes. Thus, we are interested in how different magnet types alter the ion fraction and total current. We investigated the current and ion fraction of the beam with no magnets, Sm2Co17 magnets (grade 28), and Nd2Fe14B (grade N50H). The configuration without magnets provides no confinement, and the Nd2Fe14B magnets provide the strongest confinement. We performed our measurements using the diagnostic beam line presented in Ref. [winklehner_first_2018], containing a dipole magnet and a Faraday cup with 8 mm entrance slit for mass analysis. We investigated each magnet type at a range of H2 mass flows (0.25 SCCM to 1.5 SCCM) and discharge voltages (60 V to 140 V).
In our preliminary analysis, we observed a trade-off between the total beam current and the $\mathrm{H}_{2}^{+}$ fraction: with lower confinement, the $\mathrm{H}_{2}^{+}$ fraction was high, but the current was low. This result was expected; lower confinement and lower plasma density leads to less time and fewer reactants for the ions to undergo processes that result in $\mathrm{H}_{2}^{+}$ dissociation or the pick-up of another proton (producing H+${}_{3}^{+}$). However, the lower plasma density also leads to a lower total beam current.
@@
2.1.5.5 Results with the MIST-2 back plate
Figure 2.5: A typical spectrum with the new MIST-2 back plate. The total beam current was 3.43 mA, the $\mathrm{H}_{2}^{+}$ fraction $\approx 60$%. For a discussion about the small contaminant peak, cf. text.
We already performed the first test of the MIST-2 components described in Section 2.1.2.1 by installing the new back plate (which is backward compatible) on MIST-1. We were able to run the source up to a stable beam current of 3.43 mA. We did this with the source set to 7.5 kV (the nominal source voltage of 15 kV requires a full beam current of 10 mA to form the appropriate plasma meniscus), puller set to -5 kV, lens 1 at 1.5 kV, lens 2 at -3 kV, lens 3 at 0.5 kV and the filament heating set to 150 A.
Due to the improved cooling and filament feedthroughs and holders, the source ran very stably; as we expected. A sample spectrum from this series of studies is shown in Fig. 2.5. Here, a small cluster of beam contaminants can still be seen, which we attribute to insufficient baking and out-gassing of source components, including Viton O-rings. Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [winklehner_first_2018].
@@ 2.1.6 Risks and Mitigation
Mitigation: Our current upgrade path should lead us comfortably to a total extracted current of 12 mA, particularly considering that MIST-1 has demonstrated an exceptionally low emittance, which opens up the possibility of going to even larger extraction apertures (10 mm or 12 mm). However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [alonso_isodar_2015]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$. The low ratio of 50% $\mathrm{H}_{2}^{+}$ (30 mA total current) deterred us, but as a mitigation, a VIS-type ion source paired with a dipole magnet is a viable alternative.
@@ 2.2 The Low Energy Beam Transport
Details of the LEBT design and performance were given in Refs. [waites_matching_2021, waites_low_2022, waites_high_2022] here we summarize and update these.
@@ 2.2.1 Design Overview
At this stage, ions emerge from the ion source. The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig. 2.1). These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the rest of the LEBT. A conical flange is connected to the end of the six-way cross. The conical shape minimizes beam loss on exit.
Figure 2.6: Left: Photo of Alpha Magnetics, Inc. steerers [2562NotreDame]. Right: Result of LEBT steering simulations. It can be seen that several millimeters offset can be achieved without significant change in emittance. From Ref. [waites_high_2022].
Following the six-way cross lies a diagnostic segment. A gate valve separates the six-way cross from this segment. The diagnostic section comprises a second cross, which can be used for additional vacuum attachments, diagnostic readouts, and a chopping electrode. The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning. A flange in place toward the end allows space for an AC Current Transformer (ACCT) [bergozACCT] which characterizes the beam immediately prior to entering the RFQ. The diagnostic segment is illustrated in Fig. 2.1.
@@ 2.2.2 Beam Steering & Chopping
The diagnostic segment also serves to steer the beam into the RFQ or terminate the beam as needed. Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ. In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid magnet are placed around the beamline. These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ. The optimal fields were calculated using COMSOL and beam dynamics simulated in WARP. The effect of the steerers is demonstrated in Fig. 2.6.
In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam. The beam chopper is built into the cross in the diagnostic section. The chopper comprises a deflection plate, followed by a copper ring for particle dumping press-fitted into a steel beam pipe which in turn is surrounded by a copper water cooling jacket. The deflection plate can be put at a high potential relative to the cross, which is grounded, diverting the beam such that it terminates on the copper ring. See Fig. 2.8. This method permits controlling the average beam power, without changing the charger per bunch resulting from the RFQ. This is required to maintain vortex motion in the cyclotron.
A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered. We need to ensure that the dump is capable of dissipating this amount of power. The system was modeled by assuming a water cooling line at room temperature. To be conservative, the simulation was run using four times the actual anticipated beam power, i.e. 720 W. The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper. The COMSOL simulation is shown in Fig. 2.8
Figure 2.7: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring. From Ref. [waites_high_2022].
Figure 2.8: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.From Ref. [waites_high_2022].
@@ 2.2.3 Beam Dynamics
Space charge is included in all calculations. However, as is well-known, in the absence of external electrical fields, slow secondary charged particles from the interaction of the beam with the residual gas can be accumulated inside the beam envelope, effectively neutralizing the beam charge. This process is called “space charge compensation” [winklehnerSpaceChargeCompensation2015]. In all LEBT simulations, we estimate space charge compensation according to the model presented in Ref. [winklehnerSpaceChargeCompensation2015]. It is usually above 80%, alleviating the detrimental effect of space charge in some parts of the LEBT. However, we also test our designs with no compensation and find that, while beam losses are slightly higher, we can still obtain the required beam inside the cyclotron, albeit requiring about 20% higher beam from the source.
For the IBSimu simulations a 12 mA beam composed of 80% H+${}_{2}^{+}$, 10% H+, and 10% H+${}_{3}^{+}$ was used. This provides ample current for IsoDAR assuming 50% global transmission efficiency while using current from the ion source that has already been empirically demonstrated. As illustrated in Fig. 2.9 and in Table 2.2 the LEBT can match the required input parameters for the RFQ. Moreover, as illustrated in Fig. 2.10 this match is achieved with emittance that surpasses that used for baseline RFQ simulations.
Figure 2.9: A temperature plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. [waites_high_2022]
Figure 2.10: Phase space from the LEBT simulation (left) vs phase space for ideal beam input to the RFQ (right). From Ref. [waites_high_2022].
Table 2.2: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
@@ 2.3 The Radiofrequency Quadrupole
Figure 2.11: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [waites_high_2022].
The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT is injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke. A schematic of the RFQ and cyclotron can be seen in Fig. 3.1 where only about a third of the RFQ is visible above the top surface of the cyclotron. We show a cut view of the RFQ in Fig. 2.11. Of note are the small diameter of the RFQ (27.6 cm) while operating at a low resonant frequency of 32.8 MHz. This is achieved by utilizing the so-called “split-coaxial” mode where the vanes are attached longitudinally at the entrance and exit flanges instead of radially at the mantle. The main operating parameters of the RFQ are listed in Tab. 2.3. The RFQ accelerates only moderately from 7.5 keV/amu to 35 keV/amu and acts mainly as a very efficient buncher. This leads to a conservatively estimated power consumption of $<6$ kW.
Table 2.3: A summary of the design parameters for the RFQ.

### v4

@@ 4.1.3 Electrical Design
Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig. 10. The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 standard cubic centimeter per minute (SCCM) full range. This system interfaces with the EPICS PC control system which is documented in detail in Ref. [34].
The filament is heated by an TDK Lambda (8.4 V, 300 A max) power supply. In addition to this power supply which is used to heat the filament, a second power supply (Matsusada REK) is connected to the filament and to the source chamber to maintain a potential difference of 80-140 V between them in order to facilitate electron discharge of up to 24 A max (typical during our tests was 5-10 A). The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape. Power is transferred to the platform via two isolation transformers of 2 kVA each. A schematic of the wiring of the ion source is illustrated in Fig. 11.
The ion source (platform) is lifted to high voltage using a Matsusada AU, 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU, 7.5 mA, 20 kV power supplies of either negative or positive polarity.
@@
Figure 11: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
@@ 4.1.4 Beam Dynamics
We simulated beam formation and shaping in the extraction system with the IBSimu package [37], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process. IBSimu also contains a 1D plasma sheath model which is a good approximation for the plasma-beam interface at the source aperture. Our extraction simulations of the MIST-1 ion source followed by a WARP [38] simulation of the LEBT yielded excellent agreement with measurements [31, 33]. Example trajectories of the new MIST-1 extraction system and RFQ matching LEBT are shown in Fig. 9.
An extraction system simulation of the MIST-2 system is shown in Fig. 12. The simulated normalized RMS emittance for the new MIST-2 extraction system using an 8 mm diameter plasma aperture is 0.077 $\pi$-mm-mrad. Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [31].
Figure 12: IBSimu Simulation of MIST-2 extraction. Left: particle trajectories in the z-x plane, Right: x-x’ phase space 25.5 cm after the extraction aperture. Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
@@ 4.1.5 Performance
Goals & Measurement Methods
@@
[…]
  • Low emittance: << 0.1 $\pi$-mm-mrad (RMS, normalized)
@@
Optimization of Ion Source Parameters
@@
[…]
@@
Discharge current optimization. One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 A to 8 A. No strong dependence on current was observed.
Latest Optimized Results
[…]
The main bottleneck for an increased beam current historically has been the extraction system, not the actual initial plasma generation. The extraction system characterized in this work was designed for initial testing and commissioning. We have recently replaced it with the version presented in this chapter. Our high-fidelity simulations suggest that the new extraction system will be able to transport and focus the nominal beam current the source can generate, removing this bottleneck.
Magnet Study
@@
[…]
@@
Results with the MIST-2 back plate
Figure 13: A typical spectrum with the new MIST-2 back plate. The total beam current was 3.43 mA, the $\mathrm{H}_{2}^{+}$ fraction $\approx 60$%. See main text for a discussion on the small contaminant peak.
We already performed the first test of the MIST-2 components described in Section 4.1.2 by installing the new back plate (which is backward compatible) on MIST-1. We were able to run the source up to a stable beam current of 3.43 mA. We did this with the source set to 7.5 kV (the nominal source voltage of 15 kV requires a full beam current of 10 mA to form the appropriate plasma meniscus), puller set to -5 kV, lens 1 at 1.5 kV, lens 2 at -3 kV, lens 3 at 0.5 kV and the filament heating set to 150 A.
Due to the improved cooling and filament feedthroughs and holders, the source ran very stably; as we expected. A sample spectrum from this series of studies is shown in Fig. 13. Here, a small cluster of beam contaminants can still be seen, which we attribute to insufficient baking and out-gassing of source components, including Viton O-rings. Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [31].
@@ 4.1.6 Risks and Mitigation
Mitigation: Our current upgrade path should lead us comfortably to a total extracted current of 12 mA, particularly considering that MIST-1 has demonstrated an exceptionally low emittance, which opens up the possibility of going to even larger extraction apertures (10 mm or 12 mm). However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [39]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$. The low ratio of 50% $\mathrm{H}_{2}^{+}$ (30 mA total current) deterred us, but as a mitigation, a VIS-type ion source paired with a dipole magnet is a viable alternative.
@@ 4.2 The Low Energy Beam Transport
Details of the LEBT design and performance were given in Refs. [40, 41, 36] here we summarize and update these. It should be noted that the full LEBT design for the HCHC-XX as described here is at the level of a PDR. The only exception is the system of copper electrodes shown in Fig. 9, which we already tested and reported on in the previous section.
@@ 4.2.1 Design Overview
At this stage, ions emerge from the ion source. The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig. 9). These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the rest of the LEBT. A conical flange is connected to the end of the six-way cross. The conical shape minimizes beam loss on exit.
Figure 14: Left: Photo of Alpha Magnetics, Inc. magnetic steerers [42]. Right: Result of LEBT steering simulations. It can be seen that several millimeters offset can be achieved without significant change in emittance. From Ref. [36].
Following the six-way cross lies a diagnostic segment. A gate valve separates the six-way cross from this segment. The diagnostic section comprises a second cross, which can be used for additional vacuum attachments, diagnostic readouts, and a chopping electrode. The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning. A flange in place toward the end allows space for an AC Current Transformer (ACCT) [43] which characterizes the beam immediately prior to entering the RFQ. The diagnostic segment is illustrated in Fig. 9.
@@ 4.2.2 Beam Steering & Chopping
The diagnostic segment also serves to steer the beam into the RFQ or terminate the beam as needed. Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ. In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid magnet are placed around the beamline. These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ. The optimal fields were calculated using COMSOL and beam dynamics simulated in WARP. The effect of the steerers is demonstrated in Fig. 14.
In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam. The beam chopper is built into the cross in the diagnostic section. The chopper comprises a deflection plate, followed by a copper ring for particle dumping press-fitted into a steel beam pipe which in turn is surrounded by a copper water cooling jacket. The deflection plate can be put at a high potential relative to the cross, which is grounded, diverting the beam such that it terminates on the copper ring. See Fig. 16. This method permits controlling the average beam power, without changing the charger per bunch resulting from the RFQ. This is required to maintain vortex motion in the cyclotron.
A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered. We need to ensure that the dump is capable of dissipating this amount of power. The system was modeled by assuming a water cooling line at room temperature. To be conservative, the simulation was run using four times the actual anticipated beam power, i.e., 720 W. The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper. The COMSOL simulation is shown in Fig. 16
Figure 15: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring. From Ref. [36].
Figure 16: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior. From Ref. [36].
@@ 4.2.3 Beam Dynamics
Space charge is included in all calculations. However, as is well-known, in the absence of external electrical fields, slow secondary charged particles from the interaction of the beam with the residual gas can be accumulated inside the beam envelope, effectively neutralizing the beam charge. This process is called “space charge compensation” [44]. In all LEBT simulations, we estimate space charge compensation according to the model presented in Ref. [44]. It is usually above 80%, alleviating the detrimental effect of space charge in some parts of the LEBT. However, we also test our designs with no compensation and find that, while beam losses are slightly higher, we can still obtain the required beam inside the cyclotron, albeit requiring about 20% higher beam from the source.
[…]
Figure 17: A contour plot showing the Twiss parameters required for optimal transmission through the RFQ. The indicated point shows the parameters at the end of the LEBT. From Ref. [36]
Figure 18: Phase space from the LEBT simulation (left) versus the phase space for the ideal beam input to the RFQ (right). From Ref. [36].
Table 3: A comparison of LEBT Twiss parameters and baseline RFQ parameters.
@@ 4.3 The Radiofrequency Quadrupole
Figure 19: Cut view of the IsoDAR RFQ. In the split-coaxial design, vanes are attached to the front and end plates rather than radially to the mantle. From Ref. [36].
The $\mathrm{H}_{2}^{+}$ beam coming from the LEBT will be injected into a Radio Frequency Quadrupole (RFQ) that is partially embedded in the cyclotron yoke. A schematic of the RFQ and cyclotron can be seen in Fig. 21 where only about a third of the RFQ is visible above the top surface of the cyclotron. We show a cut view of the RFQ in Fig. 19. Of note are the small diameter of the RFQ (27.6 cm) while operating at a low resonant frequency of 32.8 MHz. This is achieved by utilizing the so-called “split-coaxial” mode where the vanes are attached longitudinally at the entrance and exit flanges instead of radially at the mantle. The main operating parameters of the RFQ are listed in Table 4. The RFQ accelerates only moderately from 7.5 keV/amu to 35 keV/amu and acts mainly as a very efficient buncher. This leads to a conservatively estimated power consumption of <6<6 kW.
Table 4: A summary of the design parameters for the RFQ.